OpenStudy (anonymous):

write and equation for 3 less than 5 times another and their sum is 729

OpenStudy (saifoo.khan):

5x - 3 = 729

OpenStudy (anonymous):

thats what i thought thanks

OpenStudy (saifoo.khan):

noprob,

OpenStudy (anonymous):

5t-3=729

OpenStudy (anonymous):

but i dont feel it is correct!!!

OpenStudy (anonymous):

what would the answer be to that? cuz im supposed to get 2 answers i dont think its right either

OpenStudy (anonymous):

3 less than 5 times another and their sum is 729

OpenStudy (anonymous):

read it carefully

OpenStudy (anonymous):

it would be 5x=729 and that answer minus 3?

OpenStudy (anonymous):

Let us assume we have two numbers X and Y where X =5Y - 3 and X+Y = 729 on solving this u will get Y = 122 X = 5*122 - 3 = 610 - 3 = 607

OpenStudy (anonymous):

your are brilliant thank you!

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

wht do u do?

OpenStudy (anonymous):

for what?

OpenStudy (anonymous):

means studying or working?

OpenStudy (anonymous):

just

OpenStudy (anonymous):

homework

OpenStudy (anonymous):

k